OpenStudy (anonymous):

(1/2)x^-1/2=1 solve for x

OpenStudy (anonymous):

x=1/4

OpenStudy (anonymous):

hey can u please help me

OpenStudy (anonymous):

...

OpenStudy (anonymous):

x^-1/2 = 1/sqrt(x) your equation is (1/2)x^-1/2=1 square it on both the sides, you get (1/4)(1/x)=1 1/(4x)=1 4x=1 x=1/4

OpenStudy (anonymous):

Multiply both sides by 2:\[x^{-1/2}=2\]Take both sides to the exponent -2:\[x=2^{-2}=\frac{1}{2^2}=\frac{1}{4}\]